how do i solve 3x=2y
y=x+2

Answer by gwendolyn(128)

You have two variables and two equations, so the problem is solvable.
One of the equations we have is 'y=x+2'.
This equation has the variable 'y' alone on one side of the equation, so it can be substituted into the other equation, '3x=2y'.
When we substitute our y value into our other equation, we get
3x=2(x+2)
First, we can distribute the 2 on the right hand side of the equation.
3x=2x+4
Next, we can subtract 2x from both sides of this equation.
3x-2x=2x+4-2x
x=4
Finally, we have the equation for our 'x' variable. We can take this value and plug it in to the other equation, 'y=x+2', to find the value for our other variable.
y=x+2
y=4+2
y=6
Now, we know that x is equal to 4 and y is equal to 6.